OpenStudy (anonymous):

which expression is equivalent to 2g/3h?

OpenStudy (anonymous):

hmmmmm

OpenStudy (luckycoins888):

where are the experessions

OpenStudy (anonymous):

6^(13)/6^(6) = 6^(13-6) = 6^(7)

OpenStudy (anonymous):

*note that you can only minus one power from another when you have the same root, in this case, 6.

OpenStudy (anonymous):

14g/21h 6g/6h 2g+7/3h+7

OpenStudy (anonymous):

yes

OpenStudy (luckycoins888):

14g/21h

OpenStudy (anonymous):

yep 14g/21h

OpenStudy (anonymous):

thank you